If the position vector of one end of a chord through the focus of the parabola y2= 8x is 1/2i + 2j, find the position vector of the other end.


1
Expert's answer
2021-09-01T08:13:01-0400

Let "x_0i+y_0j" the position vector of the other end of a chord. Then "x_0=y_0^2\/8" .


The parameter of the parabola is


"2p=8\Rightarrow \dfrac{p}{2}=2"

The quadratic term in the equation is the y so the axis of the parabola coincides with the OX axis. Furthermore, the parabola lies on the positive side of the OX axis, since the coefficient accompanying the non-quadratic term (in this case the x) is 8 which is positive, so the focus lies at


"F\left( \dfrac{p}{2},0\right)=F(2,0)"

We have that:


"\dfrac{2-0}{\frac12-2}=\dfrac{y_0-0}{y_0^2/8-2}"

"2\left( \dfrac{y_0^2}{8}-2\right) =-\dfrac32y_0"

"y_0^2+6y_0-16=0\Rightarrow y_0=-8,\quad y_0=2"

The position vector of the other end is "y_0 = -8" and "x_0 = y_0^2\/8 = 8"


"\boxed{8i-8j}"

"\dfrac{}{}"